# NeuSemSlice: Towards Effective DNN Model Maintenance via Neuron-level Semantic Slicing

### v1

@@ 5.5. RQ4: Model Incremental Development
Answer to RQ4: Our method reduces the average accuracy loss for the original task to 2.79% and achieves 89.77% average accuracy in the incremental model development task. Combined with replay, it consistently excels across all settings, reducing the average accuracy loss for the original task to 0.25% and increasing average accuracy to 90.76%, effectively countering catastrophic forgetting.
@@ 5.6.1. External Threats
The threat to external validity arises from the particular settings we have chosen for our study, which means raising the generalizability concern of our proposed approach. To mitigate these threats, we have selected diverse evaluation settings. In more detail, we select four widely utilized datasets–MNIST, Fashion-MNIST, CIFAR10, and ImageNet, and employ five classical DNN architectures– MLP, AlexNet, VGG16, ResNet101, and VGG19. The experiments are conducted under six settings. This selection is intended to enhance the generalizability of our findings across common scenarios within the domain.
@@ 5.6.2. Internal Threats
Internal validity threats stem from the tools utilized in our research, which include captum (Kokhlikyan et al., 2020) and cleverhans (Papernot et al., 2018). Additionally, there is a potential threat concerning the accurate reproduction of the contribution assessment metrics and the model modularization framework DeepArc in PyTorch (Paszke et al., 2019b). The use of various deep learning frameworks, such as TensorFlow (Abadi et al., 2015), may impact the results. Additionally, the inherent randomness in model training poses a potential threat to the internal validity. We addressed this issue by repeating key experiments more than five times and reporting the mean values. Finally, as mentioned in Section 5.4, the resilience of the model’s re-adaptation to adversarial attacks requires enhancement. This issue, though orthogonal to the main focus of this paper, highlights the need for future work to tackle and reduce its impact.
@@ 6. Conclusion
This paper introduces NeuSemSlice, an innovative framework that integrates semantic slicing with semantic-aware model maintenance for deep neural networks. Semantic slicing identifies critical neurons at a granular level, categorizing and merging them across layers based on semantic similarities. This approach markedly enhances the effectiveness of these neurons in model maintenance tasks, surpassing global slicing and category-wise slicing. NeuSemSlice further innovates in semantic-aware model maintenance by offering strategies for model restructuring that preserve critical neurons, efficient model readaptation through tuning only these neurons, and incremental model development that maintains critical neurons while training non-critical ones. The effectiveness of NeuSemSlice is demonstrated through comprehensive evaluations, showing its superiority over baselines in all three aspects of model maintenance.

### v2

@@ 5.5. RQ4: Model Incremental Development
Answer to RQ4: In Experiment Scenario 1, our method reduces the average accuracy loss for the original task to 2.79% and achieves 89.77% average accuracy in the incremental model development task. Combined with replay, it consistently excels across all settings, reducing the average accuracy loss for the original task to 0.25% and increasing average accuracy to 90.76%, effectively countering catastrophic forgetting. Similar trends are also observed in Experiment Scenario 2.
Figure 6. Comparison Results of Performance for Semantic Slicing, Global Slicing, and Category-Wise Slicing on MobileNet (Left: Based on DeepLIFT, Right: Based on Taylor). Ideally, Dots (Lines) Located at the Upper Right Are Favored for Their Higher Compression Rates and Better Accuracy Performance, Which Also Suggest That the Corresponding Slicing Methods Better Identify Critical Neurons in the Model.
5.6. Applications in Software Engineering: A Case Study on MobileNet
To further demonstrate the applicability of our approach in software engineering, particularly in resource-constrained or embedded deployment scenarios, we extend our evaluation to the widely adopted MobileNet architecture (Howard et al., 2017). MobileNet is extensively utilized in real-world environments, ranging from embedded systems to intelligent IoT devices, making it a strong representative for mission-critical deployments. Specifically, we continue using CIFAR-10 as the target dataset and adopt the latest MobileNet V3 as the target model, conducting a case study under the same experimental settings. Through this evaluation, we aim to assess whether NeuSemSlice can consistently facilitate effective model maintenance in architectures that closely align with real-world software systems.
Experimental Results. Figure 6 presents the effectiveness comparison of semantic slicing, global slicing, and category-wise slicing on MobileNet using DeepLIFT and Taylor. The results clearly indicate that, at comparable compression rates (i.e., when identifying a similar proportion of critical neurons), semantic slicing more accurately identifies critical neurons in the model due to its semantic-based approach. This leads to improved model accuracy under equivalent compression rates. Specifically, for Taylor, when the model compression rate is approximately 50%, semantic slicing achieves an accuracy of 88.41%, whereas global slicing and category-wise slicing yield accuracies of 13.90% and 36.25%, respectively. DeepLIFT exhibits a similar trend. Additionally, global slicing and category-wise slicing maintain reasonable accuracy only when the compression rate is reduced to 20% or lower.
Furthermore, the three slicing methods identified using Taylor are applied in Model Restructure and integrated with existing pruning techniques to evaluate their impact on model compression efficiency. Experimental results demonstrate that for Global and ERK, Model Restructure with semantic slicing enables the model to achieve a compression rate of approximately 60% within only two pruning iterations, while maintaining an accuracy of 88.41%. In contrast, directly applying Global or ERK requires approximately nine iterations to reach a similar compression rate and model accuracy. For the Uniform+ method, without semantic slicing, the model undergoes nine pruning iterations to achieve a compression rate of 61.26%, at which point accuracy drops to 19.34%. However, when assisted by semantic slicing in Model Restructure, only two pruning iterations are needed to reach a compression rate of 61.71%, while accuracy remains at 87.95%. In contrast, global slicing and category-wise slicing exhibit inaccuracies in identifying critical neurons, leading to substantial accuracy degradation during pruning (falling below 30%), ultimately compromising model usability.
Conclusion and Outlook. Our case study on MobileNet confirms that the proposed approach can be effectively adapted to resource-constrained or embedded software systems, enabling streamlined model maintenance without compromising performance. These promising results suggest broader applicability in other software engineering scenarios. In future work, we plan to extend our methodology to more diverse real-world applications to further demonstrate its utility.
@@ 5.7.1. External Threats
The threat to external validity arises from the particular settings we have chosen for our study, which raises concerns about the generalizability of our proposed approach. To mitigate these threats, we have selected diverse evaluation settings. In more detail, we select four widely utilized datasets-MNIST, Fashion-MNIST, CIFAR10, and ImageNet, and employ six classical DNN architectures-MLP, AlexNet, MobileNet,VGG16, ResNet101, and VGG19. The experiments are conducted under seven settings. This selection is intended to enhance the generalizability of our findings across common scenarios within the domain.
@@ 5.7.2. Internal Threats
Internal validity threats stem from the tools utilized in our research, which include captum (Kokhlikyan et al., 2020) and cleverhans (Papernot et al., 2018). Additionally, there is a potential threat concerning the accurate reproduction of the contribution assessment metrics and the model modularization framework DeepArc in PyTorch (Paszke et al., 2019). The use of various deep learning frameworks, such as TensorFlow (Abadi et al., 2015), may impact the results. Additionally, the inherent randomness in model training poses a potential threat to the internal validity. We addressed this issue by repeating key experiments more than five times and reporting the mean values. Finally, as mentioned in Section 5.4, the resilience of the model’s re-adaptation to adversarial attacks requires enhancement. This issue, though orthogonal to the main focus of this paper, highlights the need for future work to tackle and reduce its impact.
@@ 6. Conclusion
This paper introduces NeuSemSlice, an innovative framework that integrates semantic slicing with semantic-aware model maintenance for deep neural networks. Semantic slicing identifies critical neurons at a granular level, categorizing and merging them across layers based on semantic similarities. This approach markedly enhances the effectiveness of these neurons in model maintenance tasks, surpassing global slicing and category-wise slicing. NeuSemSlice further innovates in semantic-aware model maintenance by offering strategies for model restructure that preserve critical neurons, efficient re-adaptation through tuning only these neurons, and incremental development that maintains critical neurons while training non-critical ones. The effectiveness of NeuSemSlice is demonstrated through comprehensive evaluations, showing its superiority over baselines in all three aspects of model maintenance.
Acknowledgments
We thank the editor and anonymous reviewers for their constructive comments. This work was supported by the National NSF of China (Grant Nos. 62302176, 62072046, 62302181), the Key R&D Program of Hubei Province (2023BAB017, 2023BAB079), and the Knowledge Innovation Program of Wuhan-Basic Research (2022010801010083).